OpenStudy (anonymous):

4 divided by (2/n)^3=

OpenStudy (anonymous):

4/(2/n)^3=4n^3/8=n^3/2

OpenStudy (anonymous):

i dont really get.it.. maybe u should put it in the equation editor and tell me how u get from each step?=)thx